OpenStudy (anonymous):

can someone help me figure out the cubic root of 3^ /1024

OpenStudy (anonymous):

You want the cube root of 1024? Please write what you want in words. It's much better.

OpenStudy (anonymous):

The is the problem on my paper 3^ / 1024 I don't have the square root symbol on my keys

OpenStudy (anonymous):

cube root of 1024 is10.079

OpenStudy (anonymous):

Sunman9 , you shouldn't write it like that. It is difficult to interpret

OpenStudy (anonymous):

but the answer according to the answer sheet is 4 3^Cubic square 16

OpenStudy (anonymous):

Just write that you want the cube root of 1024. That's it

OpenStudy (anonymous):

I have to find the root

OpenStudy (anonymous):

You have to find the nearest cube that will multiply by another number to get 1024

OpenStudy (anonymous):

I don't know where to start. I started with 512 is half of 1024

OpenStudy (anonymous):

That's very good. the cube root of 512 is 8

OpenStudy (anonymous):

so the answer is 8 cube root 2

OpenStudy (anonymous):

No Hero... According to my answer key is 4 3^cubic 16

OpenStudy (anonymous):

You're supposed to find the nearest cube. 512 is the nearest cube. The way your answer book did it is weird, but still correct

OpenStudy (anonymous):

both answers are correct

OpenStudy (anonymous):

o.k. brb

OpenStudy (anonymous):

Make sure you write you answer as \[8\sqrt[3]{2}\]

OpenStudy (anonymous):

\[\sqrt[3]{1024} \] This is the problem

OpenStudy (anonymous):

Hero

OpenStudy (anonymous):

Yes, I know, and I explained to you the answer

OpenStudy (anonymous):

I have to find the answer according to the answer key

OpenStudy (anonymous):

I never did cubic roots before so I am struggling on where to start or how to start working the problem

OpenStudy (anonymous):

\[\sqrt[3]{1024} =\sqrt[3]{512\times 2} = \sqrt[3]{512}\sqrt[3]{2} = 8\sqrt[3]{2}\]

OpenStudy (anonymous):

\[4\sqrt[3]{16}\]

OpenStudy (anonymous):

Both answers are equivalent, but mine is the most correct way to do it.

OpenStudy (anonymous):

right I did that but it did not give me the answer I need

OpenStudy (anonymous):

I'm sorry to hear that

OpenStudy (anonymous):

o.k. I appreciate the support

OpenStudy (anonymous):

This there anyone on here that is good with Finding out Cubic Root's

OpenStudy (anonymous):

You really don't get it man. My answer really is correct.